i am three times as old as my son. 5 years later,i shall be two and a half times as old as my son. how old am i and how old is my son
Answer by AAfter.Com(18)

Let my age and son's age be x and y years.
Given, x = 3y
x + 5 = 5(y + 5)/2
=> 3y + 5 = 5(y + 5)/2 = (5y + 25)/2
=> 6y + 10 = 5y + 25
=> y = 15
x = 3 X 15 = 45
Hence, my age and son's age are 45 years and 15 years.